A Food Marketing Institute found that 46% of households spend more than $125 a week on groceries. Assume the population proportion is 0.46 and a simple random sample of 76 households is selected from the population. What is the probability that the sample proportion of households spending more than S125 a week is more than than 0.38? Note: You should carefully round any z-values you calculate to 4 decimal places to match wamaps approach and calculations Answer 240(Enter your answer as a number accurate to 4 decimal places.)
Answer #1

Population proportion, p = 0.46

Sample size, n = 76

Standard error of the proportion is, se = УР(1 _ p)/n = V0.46(1-0.46)/76 = 0.0572

Probability that the sample proportions of households is more than 0.38 is,

P[p > 0.38] = P[z > (0.38 - 0.46) / 0.0572]

= P[z > -1.3986]

= 0.9190 (Rounding to 4 decimal places)